# Beyond the Hype: A dispassionate look at vision-language models in medical scenario

### v1

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities across diverse tasks, garnering significant attention in AI communities. However, their performance and reliability in specialized domains such as medicine remain insufficiently assessed. In particular, most assessments over-concentrate in evaluating VLMs based on simple Visual Question Answering (VQA) on multi-modality data, while ignoring the in-depth characteristic of LVLMs. In this study, we introduce RadVUQA, a novel Radiological Visual Understanding and Question Answering benchmark, to comprehensively evaluate existing LVLMs. RadVUQA mainly validates LVLMs across five dimensions: 1) Anatomical understanding, assessing the models’ ability to visually identify biological structures; 2) Multimodal comprehension, which involves the capability of interpreting linguistic and visual instructions to produce desired outcomes; 3) Quantitative and spatial reasoning, evaluating the models’ spatial awareness and proficiency in combining quantitative analysis with visual and linguistic information; 4) Physiological knowledge, measuring the models’ capability to comprehend functions and mechanisms of organs and systems; and 5) Robustness, which assesses the models’ capabilities against unharmonised and synthetic data. The results indicate that both generalized LVLMs and medical-specific LVLMs have critical deficiencies with weak multimodal comprehension and quantitative reasoning capabilities. Our findings reveal the large gap between existing LVLMs and clinicians, highlighting the urgent need for more robust and intelligent LVLMs. The code and dataset will be available after the acceptance of this paper.
@@ I Introduction
Certain benchmark studies have been proposed to address this issue. For instance, Lau et al. introduced VQA-RAD [1] for VQA assessment, which was built on 315 CT, MRI and Chest X-ray images with 3515 VQA pairs. SLAKE [2] evaluated LVLMs by testing on vision and knowledge-driven queries, considering more complex clinical scenarios. OmniMedVQA [3] further expanded the scale of the largest VQA dataset, with 118k images and 127k QA pairs collected from 12 modalities. These benchmarks have significantly enhanced the evaluation of medical LVLMs; however, they fall short of delivering a comprehensive analysis that fully integrates the distinct characteristics of medical imaging data. Specifically, existing studies did not systematically assess the models across critical aspects, such as quantitative and spatial reasoning, and capabilities against synthetic or unharmonised data. Instead, they merged different types of QA pairs (e.g., modality and anatomical questions) altogether and evaluated the overall evaluation score. As a result, these assessment protocols only provided limited insights, failing to elucidate specific strengths and weaknesses of different VLMs in distinct aspects. Moreover, most medical large vision-language models (LVLMs) were developed at a relatively small scale, often with around 7 billion parameters (Table I). This limited scale restricts the models’ ability to fully leverage the potential of LVLMs, which often require larger parameter counts to capture more complex patterns and nuances in data.
@@
To bridge these gaps, this study built a comprehensive dataset, RadVUQA, to evaluate LVLMs in various aspects. RadVUQA incorporates two modalities, including multi-anatomical Computed Tomography (CT) and Magnetic Resonance Imaging (MR) datasets. Unlike existing studies that primarily focus on recognizing low-level characteristics such as plane, modality, organs, and abnormalities, we built RadVUQA to assess five high-level attributes for LVLMs: 1) anatomical understanding, 2) multimodal comprehension, 3) quantitative and spatial reasoning, 4) physiological knowledge, and 5) robustness. In particular, these essential attributes were assessed based on various test settings, e.g., investigating the capabilities of LVLMs with/without a Prompt-based Chain-of-Thought (Prompt-CoT) strategy, using open-ended and close-ended questions, etc. Besides these new features as an evaluation benchmark, RadVUQA has the potential to be a VLM training set, with more complex scenarios compared with existing ones. It covers 117 and 56 organs/structures from full-body CT and MR scans, respectively, and in-depth questions to improve the model’s capabilities. In our assessment, we investigate nine solid LLVMs, including three medical-specific LVLMs (LLaVA-Med, Med-Flamingo, and RadFM), four general LVLMs (LLava, InternVL, Qwen-VL-Chat, and BLIP2). Additionally, we further include two superior commercial models, GPT-4o and Gemini-1.50-pro to test the upper-bound performance of current LVLMs. The main contribution of this paper can be summarized as
@@
  • We creatively investigate LVLMs’ capabilities in novel scenarios, e.g., the model’s performance against synthetic data and unharmonised data (varied by different imaging protocols).
@@ II-A Large Vision Language Models
Despite growing attention, there remains a lack of comprehensive evaluation of generalised LVLMs in the medical domain, especially for commercial ones (e.g., GPT4o, Gemini, etc.). To address this gap, this study conducts a thorough evaluation of highly competitive LVLMs.
@@ II-B Medical VQA benchmarks
To evaluate the capabilities of medical LVLMs, researchers have established several Medical VQA benchmarks (Table I), with VQA-RAD [1], Path-VQA [16], SLAKE[2], OmniMedVQA[3], and CAREs [17] being the most representative ones. For example, SLAKE built knowledge-driven QA pairs on a semantic-aware dataset to evaluate the LVLMs’ capabilities, and CAREs creatively evaluated more social characteristics of LVLMs such as confidence, safety, fairness, and privacy. Although efforts have been made to Medical VQA benchmarks, most studies focused on homologous aspects. In particular, recent benchmarks (e.g., CAREs and OmniMedVQA) were eager to enhance the diversity of test data (with 16 and 12 modalities, respectively) and the number of test samples, while overlooking the essence capabilities of LVLMs. For instance, although all the benchmarks assessed anatomical understanding and abnormalities, few of them investigated quantitative reasoning and spatial reasoning, while none of them explored multimodal comprehension capability. Interestingly, the common conclusion indicates that general LVLMs significantly outperform their medical-specific counterparts in terms of overall capabilities. This finding highlights the need for a deeper investigation into the specific strengths and weaknesses of LVLMs in the medical domain.
In this study, we further investigate the strengths and weaknesses of generalized LVLMs and their medical counterparts by assessing their in-depth properties when dealing with medical queries. We creatively introduce the multimodal comprehensive assessment, which evaluates the capability of LVLMs given the prior knowledge from both linguistic instructions (text prompts) and visual guidance (visual box prompts). Additionally, we carefully designed multiple-choice questions to better estimate the spatial reasoning of various LVLMs, which enables us to explore whether the LVLMs have learned knowledge in a systematic, clinicians-like manner or are merely relying on representation clipping.
@@ III-A1 Data Resources
RadVUQA-OOD was collected from multiple resources, including 250 synthetic 2D chest CT images from [20], 250 real 2D chest CT images from [21], and 63 2D animal CT scans from embodi3D [^1] https://www.embodi3d.com/. Additionally, the real images (from [21]) were augmented to simulate unharmonised data by introducing motion-blur, window-shift, noisy, sharpness, and low-resolution variants, with 1250 images (250 for each) in total.
@@ III-B1 RadVUQA-CT and RadVUQA-MRI
Basic prompt: This image is a {imaging modality} scan with human organs and structures, please answer the following question(s). Advanced prompt: This image is a {imaging modality} scan of the human {abdominal cavity}. In this image, structures such as the {small intestine}, and {left kidney} can be observed. Please answer the following question(s).
@@ III-B2 RadVUQA-OOD
Initially, CT and MRI heavily rely on image preprocessing techniques. For instance, the influence of acquisition protocols (such as different reconstruction kernels, normalisation strategies, and patient positioning) on model performance has not been thoroughly investigated. These factors are critical in real-world clinical environments where imaging conditions are far from standardized and harmonised. Guided by [22], the OOD subsets were designed to assess the models’ capability against these different scenarios, including noise, diverse contrast, sharpness, motion blur, low-dose scanning, etc. Specifically, we assess LVLMs in the following aspects:
Robustness simulate various scenarios including motion blur, biased imaging protocol (different reconstruction kernels with different contrast and sharpness), low-resolution scanning, and noisy data. For fair comparisons, the QA pair for the unharmonised data remains the same as those for OEQ1.1, OEQ1.2, and CEQ1.
@@
OEQ5. What is shown in this image? Is it a human CT scan? CEQ5. This image is a {CT scan} of the {human thoracic cavity}, it could be a synthesised one or a real one. Please choose one option below that you believe is correct.
A. It is synthesised. B. It is a real one. C. I don’t know.
@@
In general, RadVUQA comprises 10,759 images and 193,662 question-answer pairs (different context prompts combined with OEQs/CEQs), focusing on evaluating five fundamental characteristics of LVLMs. The dataset includes a diverse range of anatomical structures, with 117 from CT scans and 56 from MRI scans. Despite the fewer modalities and number of images compared to OmniMedVQA [3], RadVUQA comprises more QA pairs due to its in-depth and thorough QA framework, ensuring data diversity through its extensive coverage of anatomical structures. We want to emphasize that although anatomical attributes may seem intuitive and straightforward, they are essential for evaluating the core capabilities of LVLMs in the medical domain. These attributes are critical for assessing whether these models possess a genuine understanding of human physiological knowledge.
@@ IV Experimental settings and metrics
Multiple-Choice Accuracy: {Predicted scores normalised to [0, 1]}.
@@ V Results
RadVUQA-CT. Table II demonstrates the capabilities of different LVLMs on CT data. Specifically, Gemini-Flash achieved the highest score in anatomic understanding, with a 3.13 Response Accuracy (RA) score and a 0.68 Multiple-Choice Accuracy (MCA) score. GPT-4o outperformed all comparisons in openly answering vision & language comprehension questions, achieving a leading RA score of 1.79. For close-ended questions, Huatuo-34b outperformed GPT-4o with the MCA score of 0.55. In quantitative reasoning, GPT-4o obtained the top RA score of 2.54 and the best MCA score (0.60). Interestingly, a similar trend is witnessed in spatial reasoning, with GPT-4o achieving the best RA score (3.32) in open-ended QA and Claude-Sonnet’s 0.48 MCA score in close-ended QA. As for physiological knowledge, Huatuo-34b led with an RA score of 2.08, outperforming other LVLMs.
@@
Model Noisy Contrast Sharpness Motion Blur Low Resolution
OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination
LLaVA-Med 1.52 0.302 19.34% 1.59 0.332 20.99% 1.62 0.343 19.34% 1.6 0.282 22.10% 1.64 0.362 22.10%
Huatuo-7b 2.59 0.435 16.02% 2.59 0.440 21.55% 2.67 0.428 18.89% 2.76 0.432 15% 2.7 0.442 17.13%
Huatuo-34b 2.58 0.377 12.15% 2.56 0.410 19.34% 2.61 0.397 13.26% 2.67 0.373 14.92% 2.6 0.372 17.68%
LLaVA 2.27 0.435 57.46% 2.36 0.422 51.38% 2.29 0.452 56.91% 2.48 0.442 48.07% 2.36 0.440 50.83%
InternVL 2.83 0.430 18.99% 2.85 0.453 17.13% 2.82 0.455 20.99% 2.84 0.490 18.23% 2.84 0.467 18.33%
MiniCPM 2.41 0.445 17.13% 2.6 0.432 17.22% 2.55 0.463 12.15% 2.54 0.467 25.41% 2.45 0.438 23.20%
BLIP2 1.04 0.187 2.76% 1.04 0.168 8.83% 1.03 0.207 3.31% 1.44 0.157 10.50% 1.51 0.190 4.97%
GPT-4o 2.92 0.608 10.50% 2.96 0.615 6.67% 2.97 0.607 5.52% 3.04 0.583 5.52% 3.02 0.618 7.18%
Gemini-Flash 2.67 0.743 27.78% 2.76 0.723 27.07% 2.64 0.738 23.33% 2.75 0.733 18.78% 2.82 0.738 24.44%
Claude 2.78 0.538 11.60% 2.86 0.447 12.71% 2.8 0.500 14.92% 2.79 0.498 12.71% 2.79 0.478 8.38%
RadVUQA-OOD. Table III presents the capabilities of LVLMs on the unharmonised data. In particular, GPT-4o achieved the highest RA scores in open-ended questions under all five scenarios, with 2.92, 2.96, 2.97, 3.04, and 3.02 on noisy, contrast, sharpness, motion blur, and low-resolution data, respectively). Meanwhile, Gemini-Flash obtained the best MCA scores in all close-ended questions (with 0.743, 0.723, 0.738, 0.733, and 0.738, respectively). In addition to GPT-4o and Gemini, InternVL also achieved competitive performance with around 2.83 RA scores. Notably, GPT-4o and BLIP2 were observed with the lowest hallucination scores for the unharmonised data.

### v2

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities across diverse tasks, garnering significant attention in AI communities. However, their performance and reliability in specialized domains such as medicine remain insufficiently assessed. In particular, most assessments over-concentrate on evaluating VLMs based on simple Visual Question Answering (VQA) on multi-modality data, while ignoring the in-depth characteristics of LVLMs. In this study, we introduce RadVUQA, a novel Radiological Visual Understanding and Question Answering benchmark, to comprehensively evaluate existing LVLMs. RadVUQA mainly validates LVLMs across five dimensions: 1) Anatomical understanding, assessing the models’ ability to visually identify biological structures; 2) Multimodal comprehension, which involves the capability of interpreting linguistic and visual instructions to produce desired outcomes; 3) Quantitative and spatial reasoning, evaluating the models’ spatial awareness and proficiency in combining quantitative analysis with visual and linguistic information; 4) Physiological knowledge, measuring the models’ capability to comprehend functions and mechanisms of organs and systems; and 5) Robustness, which assesses the models’ capabilities against unharmonized and synthetic data. The results indicate that both generalized LVLMs and medical-specific LVLMs have critical deficiencies with weak multimodal comprehension and quantitative reasoning capabilities. Our findings reveal the large gap between existing LVLMs and clinicians, highlighting the urgent need for more robust and intelligent LVLMs. The code is available at https://github.com/Nandayang/RadVUQA.
@@ I Introduction
Certain benchmark studies have been proposed to address this issue. For instance, Lau et al. introduced VQA-RAD [1] for VQA assessment, which was built on 315 CT, MRI, and Chest X-ray images with 3515 VQA pairs. SLAKE [2] evaluated LVLMs by testing on vision and knowledge-driven queries, considering more complex clinical scenarios. OmniMedVQA [3] further expanded the scale of the largest VQA dataset, with 118k images and 127k QA pairs collected from 12 modalities. These benchmarks have significantly enhanced the evaluation of medical LVLMs; however, they fall short of delivering a comprehensive analysis that fully integrates the distinct characteristics of medical imaging data. Specifically, existing studies did not systematically assess the models across critical aspects, such as quantitative and spatial reasoning, and capabilities against synthetic or unharmonized data. Instead, they merged different types of QA pairs (e.g., modality and anatomical questions) altogether and evaluated the overall evaluation score. As a result, these assessment protocols only provided limited insights, failing to elucidate specific strengths and weaknesses of different VLMs in distinct aspects. Moreover, most medical large vision-language models (LVLMs) were developed at a relatively small scale, often with around 7 billion parameters (Table I). This limited scale restricts the models’ ability to fully leverage the potential of LVLMs, which often require larger parameter counts to capture more complex patterns and nuances in data.
@@
To bridge these gaps, this study built a comprehensive dataset, RadVUQA, to evaluate LVLMs in various aspects. RadVUQA incorporates two modalities, including multi-anatomical Computed Tomography (CT) and Magnetic Resonance Imaging (MR) datasets. Unlike existing studies that primarily focus on recognizing low-level characteristics such as plane, modality, organs, and abnormalities, we built RadVUQA to assess five high-level attributes for LVLMs: 1) anatomical understanding, 2) multimodal comprehension, 3) quantitative and spatial reasoning, 4) physiological knowledge, and 5) robustness. In particular, these essential attributes were assessed based on various test settings, e.g., investigating the capabilities of LVLMs with/without a Prompt-based Chain-of-Thought (Prompt-CoT) strategy, using open-ended and close-ended questions, etc. Besides these new features as an evaluation benchmark, RadVUQA has the potential to be a VLM training set, with more complex scenarios compared with existing ones. It covers 117 and 56 organs/structures from full-body CT and MR scans, respectively, and in-depth questions to improve the model’s capabilities. In our assessment, we investigate nine solid LVLMs, including three medical-specific LVLMs (LLaVA-Med and Huatuo) and four general LVLMs (LLava, InternVL, MiniCPM, and BLIP2). Additionally, we further include two superior commercial models, GPT-4o and Gemini-1.50-pro to test the upper-bound performance of current LVLMs. The main contribution of this paper can be summarized as
@@
  • We creatively investigate LVLMs’ capabilities in novel scenarios, e.g., the model’s performance against synthetic data and unharmonized data (varied by different imaging protocols).
@@ II-A Large Vision Language Models
Despite growing attention, there remains a lack of comprehensive evaluation of generalized LVLMs in the medical domain, especially for commercial ones (e.g., GPT4o, Gemini, etc.). To address this gap, this study conducts a thorough evaluation of highly competitive LVLMs.
@@ II-B Medical VQA benchmarks
To evaluate the capabilities of medical LVLMs, researchers have established several Medical VQA benchmarks (Table I), with VQA-RAD [1], Path-VQA [16], SLAKE[2], OmniMedVQA[3], and CAREs [17] being the most representative ones. For example, SLAKE built knowledge-driven QA pairs on a semantic-aware dataset to evaluate the LVLMs’ capabilities, and CAREs creatively evaluated more social characteristics of LVLMs such as confidence, safety, fairness, and privacy. Although efforts have been made to Medical VQA benchmarks, most studies focused on homologous aspects. In particular, recent benchmarks (e.g., CAREs and OmniMedVQA) were eager to enhance the diversity of test data (with 16 and 12 modalities, respectively) and the number of test samples, while overlooking the essential capabilities of LVLMs. For instance, although all the benchmarks assessed anatomical understanding and abnormalities, few of them investigated quantitative reasoning and spatial reasoning, while none of them explored multimodal comprehension capability. Interestingly, the common conclusion indicates that general LVLMs significantly outperform their medical-specific counterparts in terms of overall capabilities. This finding highlights the need for a deeper investigation into the specific strengths and weaknesses of LVLMs in the medical domain.
In this study, we further investigate the strengths and weaknesses of generalized LVLMs and their medical counterparts by assessing their in-depth properties when dealing with medical queries. We creatively introduce the multimodal comprehensive assessment, which evaluates the capability of LVLMs given the prior knowledge from both linguistic instructions (text prompts) and visual guidance (visual box prompts). Additionally, we carefully designed multiple-choice questions to better estimate the spatial reasoning of various LVLMs, which enables us to explore whether the LVLMs have learned knowledge in a systematic, clinician-like manner or are merely relying on representation clipping.
@@ III-A1 Data Resources
RadVUQA-OOD was collected from multiple resources, including 250 synthetic 2D chest CT images from [20], 250 real 2D chest CT images from [21], and 63 2D animal CT scans from embodi3D [^1] https://www.embodi3d.com/. Additionally, the real images (from [21]) were augmented to simulate unharmonized data by introducing motion-blur, window-shift, noisy, sharpness, and low-resolution variants, with 1250 images (250 for each) in total.
@@ III-B1 RadVUQA-CT and RadVUQA-MRI
Basic prompt: This image is a {imaging modality} scan with human organs and structures; please answer the following question(s). Advanced prompt: This image is a {imaging modality} scan of the human {abdominal cavity}. In this image, structures such as the {small intestine}, and {left kidney} can be observed. Please answer the following question(s).
@@ III-B2 RadVUQA-OOD
Initially, CT and MRI heavily rely on image preprocessing techniques. For instance, the influence of acquisition protocols (such as different reconstruction kernels, normalization strategies, and patient positioning) on model performance has not been thoroughly investigated. These factors are critical in real-world clinical environments where imaging conditions are far from standardized and harmonized. Guided by [22], the OOD subsets were designed to assess the models’ capability against these different scenarios, including noise, diverse contrast, sharpness, motion blur, low-dose scanning, etc. Specifically, we assess LVLMs in the following aspects:
Robustness simulate various unharmonized scenarios including motion blur, biased imaging protocol (different reconstruction kernels with different contrast and sharpness), low-resolution scanning, and noisy data. We simulated these unharmonized conditions by randomly applying multiple distortion factors to the original test data. In particular, we either applied random Gaussian noise (mean = 0, variance = 5), a motion blur (kernel sizes range from 3 to 10), downsampling with a scale factor between 0.5 and 0.9, adjusted image contrast using gamma values in the 0.8 to 1.3 range, or histogram equalization process. For fair comparisons, the QA pair for the unharmonized data remains the same as those for OEQ1.1, OEQ1.2, and CEQ1.
@@
OEQ5. What is shown in this image? Is it a human CT scan? CEQ5. This image is a {CT scan} of the {human thoracic cavity}, it could be a synthesized one or a real one. Please choose one option below that you believe is correct.
A. It is synthesized. B. It is a real one. C. I don’t know.
@@
In general, RadVUQA comprises 10,759 images and 193,662 question-answer pairs (different context prompts combined with OEQs/CEQs), focusing on evaluating five fundamental characteristics of LVLMs. The dataset includes a diverse range of anatomical structures, with 117 from CT scans and 56 from MRI scans. Despite the fewer modalities and number of images compared to OmniMedVQA [3], RadVUQA comprises more QA pairs due to its in-depth QA framework, ensuring data diversity through its extensive coverage of anatomical structures. We want to emphasize that although anatomical attributes may seem intuitive and straightforward, they are essential for evaluating the core capabilities of LVLMs in the medical domain. These attributes are critical for assessing whether these models possess a genuine understanding of human physiological knowledge.
@@ IV Experimental settings and metrics
Multiple-Choice Accuracy: {Predicted scores normalized to [0, 1]}.
@@ V Results
RadVUQA-CT. Table II demonstrates the capabilities of different LVLMs on CT data. Specifically, Gemini-Flash achieved the highest score in anatomic understanding, with a 3.13 Response Accuracy (RA) score and a 0.68 Multiple-Choice Accuracy (MCA) score. GPT-4o outperformed all comparisons in openly answering vision & language comprehension questions, achieving a leading RA score of 1.79. For close-ended questions, Huatuo-34b outperformed GPT-4o with an MCA score of 0.55. In quantitative reasoning, GPT-4o obtained the top RA score of 2.54 and the best MCA score (0.60). Interestingly, a similar trend is witnessed in spatial reasoning, with GPT-4o achieving the best RA score (3.32) in open-ended QA and Claude-Sonnet’s 0.48 MCA score in close-ended QA. As for physiological knowledge, Huatuo-34b led with an RA score of 2.08, outperforming other LVLMs.
@@
Model Noisy Contrast Sharpness Motion Blur Low Resolution Histo Equalization
OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination OEQ CEQ Hallucination
LLaVA-Med 1.52 0.302 19.34% 1.59 0.332 20.99% 1.62 0.343 19.34% 1.6 0.282 22.10% 1.64 0.362 22.10% 1.63 0.621 14.43%
Huatuo-7b 2.59 0.435 16.02% 2.59 0.440 21.55% 2.67 0.428 18.89% 2.76 0.432 15% 2.7 0.442 17.13% 2.52 0.682 13.3%
Huatuo-34b 2.58 0.377 12.15% 2.56 0.410 19.34% 2.61 0.397 13.26% 2.67 0.373 14.92% 2.6 0.372 17.68% 2.48 0.665 11.7%
LLaVA 2.27 0.435 57.46% 2.36 0.422 51.38% 2.29 0.452 56.91% 2.48 0.442 48.07% 2.36 0.440 50.83% 2.23 0.665 19.68%
InternVL 2.83 0.430 18.99% 2.85 0.453 17.13% 2.82 0.455 20.99% 2.84 0.490 18.23% 2.84 0.467 18.33% 2.56 0.663 2.13%
MiniCPM 2.41 0.445 17.13% 2.6 0.432 17.22% 2.55 0.463 12.15% 2.54 0.467 25.41% 2.45 0.438 23.20% 2.43 0.633 10.16%
BLIP2 1.04 0.187 2.76% 1.04 0.168 8.83% 1.03 0.207 3.31% 1.44 0.157 10.50% 1.51 0.190 4.97% 1.93 0.475 20.86%
GPT-4o 2.92 0.608 10.50% 2.96 0.615 6.67% 2.97 0.607 5.52% 3.04 0.583 5.52% 3.02 0.618 7.18% 2.80 0.752 3.72%
Gemini-Flash 2.67 0.743 27.78% 2.76 0.723 27.07% 2.64 0.738 23.33% 2.75 0.733 18.78% 2.82 0.738 24.44% 2.78 0.784 9.63%
Claude 2.78 0.538 11.60% 2.86 0.447 12.71% 2.8 0.500 14.92% 2.79 0.498 12.71% 2.79 0.478 8.38% 2.62 0.709 8.51%
RadVUQA-OOD. Table III presents the capabilities of LVLMs on the unharmonized data. In particular, GPT-4o achieved the highest RA scores in open-ended questions under all six scenarios, with 2.92, 2.96, 2.97, 3.04, 3.02, and 2.8 on noisy, contrast, sharpness, motion blur, low-resolution, and histogram equalization data respectively). Meanwhile, Gemini-Flash obtained the best MCA scores in all close-ended questions (with 0.743, 0.723, 0.738, 0.733, 0.738, and 0.784, respectively). In addition to GPT-4o and Gemini, InternVL also achieved competitive performance with around 2.83 RA scores. Notably, GPT-4o and BLIP2 were observed to have the lowest hallucination scores for the unharmonized data.